OpenStudy (anonymous):

limit as x approaches negative infinity 3x/(4x^2+10)^(1/2) why is this -3/2 instead of 3/2 if you end up with 3/(2+10^(1/2)/x)

OpenStudy (anonymous):

why would you do that?

OpenStudy (anonymous):

I thought you just divided by biggest x^n

OpenStudy (zarkon):

\[\sqrt{4x^2+10}=\sqrt{x^2}\sqrt{4+\frac{10}{x^2}}=|x|\sqrt{4+\frac{10}{x^2}}\]

OpenStudy (zarkon):

then for \(x<0\) we have \[\frac{x}{|x|}=-1\]

OpenStudy (turingtest):

dang, this is trickier than I anticipated I like Zarkon's way but I'm, trying to find another

OpenStudy (anonymous):

gahh i hate theory, is there any other way to explain it having to do that would be a huge pain in the retriceon the test I'm going to have to remember that specifically. I didn't even know you could relate (x^2)^(1/2) to abs(x). How is that I get a 100% on the first test to not understanding what is going on at all...

OpenStudy (turingtest):

actually the definition of \[|x|=\sqrt{x^2}\]that is a good thing to know in general

OpenStudy (turingtest):

latex question: how do you write the defined as symbol?

OpenStudy (zarkon):

this? \[\equiv\]

OpenStudy (zarkon):

I usuallt just use := so \[|x|:=\sqrt{x^2}\]

OpenStudy (anonymous):

I wish they would have taught us this earlier.... lol

OpenStudy (zarkon):

usually it is written as \[|x|=\left\{\begin{matrix}x & \text{if}& x\ge0 \\-x & \text{if}&x<0\end{matrix}\right.\]

OpenStudy (turingtest):

I mine wrong? I know they use it for some proofs

OpenStudy (turingtest):

is*

OpenStudy (zarkon):

\[|x|=\sqrt{x^2}\] is a valid definition